OpenStudy (anonymous):

please help me with this make c the subject of the formular in the following equation F = 9/5C+32 please i need step by step solution on that.

OpenStudy (anonymous):

looks like you need to solve for \(C\), i.e. isolate \(C\) on one side of the equation you have: \[F=\frac{9}{5}C+32\] we can start by substracting \(32\) to both sides, we get: \[F-32=\frac{9}{5}C\] what do you think is the next step ??

OpenStudy (anonymous):

you cross multilpy both sides and we get: C= 5/9(f-32) thanks bro for the insight

OpenStudy (anonymous):

i see you got the idea... glad to help!